OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot

OpenStudy (anonymous):

 Area of circle is Pi * r * r so 206 = pi * r^2 so radius = \[\sqrt{206/pi}\]

OpenStudy (anonymous):

= 8 feet.